OpenStudy (anonymous):

Somebody please help! I have no idea how to do this! Give an example of a consistent dependent system. Explain why the system is a consistent dependent system.

OpenStudy (anonymous):

humm........................................................

OpenStudy (anonymous):

do you know what a dependent system is?

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

ok... so a dependent system has infinite numbers.

OpenStudy (anonymous):

i mean solutions

OpenStudy (anonymous):

I don't get it. What is it exactly?

OpenStudy (anonymous):

like here is a dependent system... y=x+1 2y=2x+1

OpenStudy (anonymous):

if you simplify it, they would be identical.

OpenStudy (anonymous):

Okay, I see. But is a consistent dependent system the same thing?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

a consistent equation has at least ONE solution

OpenStudy (anonymous):

So I could just put an example like yours and it would be correct? Okay then. Thanks for the help.

Directrix (directrix):

If a system has at least one solution, it is said to be consistent . If a consistent system has an infinite number of solutions, it is dependent . When you graph the equations, both equations represent the same line.

Directrix (directrix):

x + y = 1 2x + 2y = 2 ======== The same line.

Directrix (directrix):

Because the two equations are of the same line, the graphs coincide at every point. Hence, infinitely many solutions.

OpenStudy (anonymous):

Ohhhhh! :) Thanks both of you for your help. :) Now I get it.

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

what he said.